# “Haet Bhasha aur Diskrimineshun”: Phonetic Perturbations in Code-Mixed Hinglish to Red-Team LLMs

### v1

@@ Abstract
Large Language Models (LLMs) have become increasingly powerful, with multilingual and multimodal capabilities improving by the day. These models are being evaluated through audits, alignment studies and red-teaming efforts to expose model vulnerabilities towards generating harmful, biased and unfair content. Existing red-teaming efforts have previously focused on the English language, using fixed template-based attacks; thus, models continue to be susceptible to multilingual jailbreaking strategies, especially in the multimodal context. In this study, we introduce a novel strategy that leverages code-mixing and phonetic perturbations to jailbreak LLMs for both text and image generation tasks. We also introduce two new jailbreak strategies that show higher effectiveness than baseline strategies. Our work presents a method to effectively bypass safety filters in LLMs while maintaining interpretability by applying phonetic misspellings to sensitive words in code-mixed prompts. Our novel prompts achieve a 99% Attack Success Rate for text generation and 78% for image generation, with Attack Relevance Rate of 100% for text generation and 95% for image generation when using the phonetically perturbed code-mixed prompts. Our interpretability experiments reveal that phonetic perturbations impact word tokenization, leading to jailbreak success. Our study motivates increasing the focus towards more generalizable safety alignment for multilingual multimodal models, especially in real-world settings wherein prompts can have misspelt words.
@@ 2 Related Work
Code-mixing (CM) increases the performance Shankar et al. (2024) and capabilities Zhang et al. (2024) of LLMs in multilingual settings. However, multilingual safety evaluation of LLMs remains underexplored Song et al. (2024). In this study, we study code-mixing with phonetic perturbations of sensitive words as an attack vector (as opposed to code-switching Yoo et al. (2024)) to evaluate safety alignment of LLMs. This novel jailbreak strategy successfully jailbreaks even SOTA models like Llama 3 and ChatGPT 4o-mini.
@@ 3.1 Datasets Benchmarked
Prompts for text generation: We evaluate our red-teaming strategy by modifying data from three benchmark datasets that have prompts for studying model vulnerabilities, refusal training and compliance with harmful queries [^3] See Appendix for dataset descriptionsHarmfulQA Bhardwaj and Poria (2023), NicheHazardQA Hazra et al. (2024) and TechHazardQA Banerjee et al. (2024). We randomly sample 20 prompts from each category in each dataset, yielding a total of 460 prompts across 23 categories. All prompts are originally in the English language. We attempted to generate code-mixed prompts with phonetic perturbations using automated means Le et al. (2022) but the results were unsatisfactory. Thus, we sample subsets from each dataset and manually generate code-mixed prompts with phonetic perturbations in the Hindi and English languages, described later.
@@ 3.2 Models Evaluated
In this study, we benchmark four instruction-tuned LLMs of comparable sizes $\approx$ 8B parameters, with varying levels of multilingual capabilities [^4] See Appendix for model descriptionsChatGPT-4o-mini Hurst et al. (2024), Llama-3-8B-Instruct Dubey et al. (2024), Gemma-1.1-7b-it Team et al. (2024), Mistral-7B-Instruct-v0.3 Jiang et al. (2023). For the multimodal image generation task, we use ChatGPT-4o-mini Hurst et al. (2024).
@@ 3.3 Jailbreak Templates
In this study, we benchmark three existing jailbreak templates on the LLMs [^5] More details are shared in the Appendix.Opposite Mode (OM), AntiLM, AIM Shen et al. (2024b) using code-mixed and phonetically perturbed inputs. For the image generation task, we test with a Base template– instructing image generation without asking for clarifications for generation style. We also devise two new jailbreaking templates
@@ 4.1 Generating Input Prompts
We use a 3-step process to generate the code-mixed prompts and phonetic perturbations for both text and image generation tasks [^6] See Appendix for an example walkthrough..
@@ 4.3 Interpreting Phonetic Perturbations
  • We select a small subset of the dataset, specifically with $AASR_{CM}\leq 0.33$, $AASR_{CMP}\geq 0.5$ while ensuring that $AARR_{CMP}\geq AARR_{CM}$.
@@ 5 Results & Observations
In Table 1 we report the average Attack Success Rate (AASR) for all datasets, prompts, models and jailbreak templates. We first note that Gemma and Mistral report the highest AASR across jailbreak templates and input prompts. In fact, if we do not use any jailbreaking template, attacks still succeed at least 65% of the time on Gemma and 68% on Mistral. This baseline experiment shows that models are still vulnerable to classic red-teaming attacks despite evidence of harm, and using code-mixed (CM) or phonetically perturbed (CMP) prompts in fact increases the AASR, thus showing the effectiveness of our approach. ChatGPT and Llama are fairly robust to the different types of input, irrespective of the jailbreaking template being used. We also note that combining CM prompts with any of the jailbreak templates again results in $\simeq$ 0 AASR for ChatGPT and Llama, showing that their safety filters put precedence on template based attacks rather than the language of the input; Gemma and Mistral still maintain high AASR across the templates. We also note that our proposed jailbreak technique– Sandbox performs the best amongst all, for both ChatGPT and Llama. This could be because the models are not aware of this template. We can conclude that two out of the four models are successfully red-teamed using our input format and jailbreak template. Next, we attempt to understand how relevant the outputs are for each input prompt.
@@
In Figures 35, we plot the sequence attribution scores using the method of Integrated Gradients Sundararajan et al. (2017), for a given input prompt in the English language, its code-mixed (CM) variant and its code-mixed variant with phonetic perturbations (CMP) for the Llama-3-8B-Instruct LLM. Recall that we study the attribution scores for generating safe outputs; thus, if a token has a high score, it implies that the token contributes more towards triggering the safety filter and if it scores low, then it has a lower impact towards the generated output. In Figure 3 we observe how the tokens– “hate”, “speech” and “discrimination” have a high attribution score at the embedding layer, 1st decoder layer and the 8th decoder layer. This means that these tokens are primarily responsible for generating the safe response from the LLM. “hate” and “speech” retain their high importance even at the 16th decoder layer. A similar observation can be made in Figure 4, for the same prompt written in a code-mixed manner, where “hate” and “speech” are written in the English language. Thus, standard code-mixing may not be enough to bypass the safety filters of LLMs. On the other hand, in Figure 5, we note that the prompt, after phonetic perturbations, generates radically different tokens– “hate” $\rightarrow$ “haet” tokenized as “ha” + “et” and, “discrimination” $\rightarrow$ “bhed bhav” tokenized as “b” + “hed” + “b” + “ha” + “av”. This leads to a change in the attribution scores as well, which are now low for the sensitive words and thus do not trigger the safety filters. A similar observation can be made for other prompts. Thus, based on our observations, we can state that phonetic perturbations lead to input tokenisation in a way that impacts the safety filters of LLMs, thus allowing attackers to generate harmful outputs.
@@ 6 Discussion
  • Improved safety measures for all models: While Llama and ChatGPT show resistance to standard jailbreak attacks, Gemma and Mistral are easily jailbroken, even in English. This highlights the need for better evaluation and alignment methods for these models, especially considering their wide-scale availability. In multilingual and multimodal settings, especially code-mixed settings, safety filters of all models degrade drastically, revealing the need for better efforts in multilingual multimodal red teaming and alignment.
  • Template-based safety measures do not generalize well: Guardrails of models that are safety trained against template based attacks do not generalize well to attacks that deviate from set patterns, by including elements such as perturbations or code-mixing. This highlights the need for more generalizable safety training methods that are robust to deviations from standard attack patterns.
  • Input tokenization determines safe outputs: If input tokens are out of vocabulary or perturbed, the safety filters get bypassed easily, but the model is able to interpret the instructions in the input prompt and generate harmful responses. Thus a more robust tokenization strategy needs to be developed.
@@ 7 Conclusion
Future Work: An immediate area of future work is to align the models based on the findings from the explainability experiments. We also plan to expand our red-teaming approach to other output modalities like speech. We plan to scale our efforts to more prompt categories, models and jailbreak templates.
@@ A.1.3 Jailbreaking Templates
In all cases, the templates contain instructions for allowing the jailbroken counterparts of the bots to bypass any safety guardrails, and not inform the user about legal and ethical guidelines The jailbreak templates are used as system prompts for all the models except for Gemma-1.1-7b-it which does not allow using system prompts. For Gemma, we append it to the main prompt as a prefix.

### v2

@@ Abstract
Recently released LLMs have strong multilingual & multimodal capabilities. Model vulnerabilities are exposed using audits and red-teaming efforts. Existing efforts have focused primarily on the English language; thus, models continue to be susceptible to multilingual jailbreaking strategies, especially for multimodal contexts. In this study, we introduce a novel strategy that leverages code-mixing and phonetic perturbations to jailbreak LLMs for both text and image generation tasks. We also introduce two new jailbreak strategies that show higher effectiveness than baselines. Our work presents a method to effectively bypass safety filters in LLMs while maintaining interpretability by applying phonetic misspellings to sensitive words in code-mixed prompts. We achieve a 99% Attack Success Rate for text generation and 78% for image generation, with Attack Relevance Rate of 100% for text generation and 95% for image generation for the phonetically perturbed code-mixed prompts. Our interpretability experiments reveal that phonetic perturbations impact word tokenization, leading to jailbreak success. Our study motivates increasing the focus towards more generalizable safety alignment for multilingual multimodal models, especially in real-world settings wherein prompts can have misspelt words.
@@ 2 Related Work
Code-mixing (CM) increases the performance Shankar et al. (2024) and capabilities Zhang et al. (2024) of LLMs in multilingual settings. Prior works in multilingual red-teaming of LLMs involves evaluation Shen et al. (2024a), jailbreaking Deng et al. (2023) as well as alignment Song et al. (2024) strategies. Unlike code-switching Yoo et al. (2024) which directly inserts words from other languages in their native scripts, code-mixing uses the script of the primary language to insert words from other languages, making it a more natural communication style as observed in SMS and internet conversations Thakur (2021). In this study, we study code-mixing with phonetic perturbations of sensitive words as an attack vector to evaluate safety alignment of LLMs. This novel jailbreak strategy successfully jailbreaks even SOTA models like Llama 3 and ChatGPT 4o-mini.
@@ 3.1 Datasets Benchmarked
Prompts for text generation: We evaluate our red-teaming strategy by modifying data from three benchmark datasets that have prompts for studying model vulnerabilities, refusal training and compliance with harmful queries [^3] See Appendix for more details.HarmfulQA Bhardwaj and Poria (2023), NicheHazardQA Hazra et al. (2024) and TechHazardQA Banerjee et al. (2024). We randomly sample 20 prompts from each category in each dataset, yielding a total of 460 prompts across 23 categories. All prompts are originally in the English language. We attempted to generate code-mixed prompts with phonetic perturbations using automated means Le et al. (2022) but the results were unsatisfactory. Thus, we sample subsets from each dataset and manually generate code-mixed prompts with phonetic perturbations in the Hindi and English languages, described later.
@@ 3.2 Models Evaluated
In this study, we benchmark four instruction-tuned LLMs of comparable sizes $\approx$ 8B parameters, with varying levels of multilingual capabilities [^3] See Appendix for more details.ChatGPT-4o-mini Hurst et al. (2024), Llama-3-8B-Instruct Dubey et al. (2024), Gemma-1.1-7b-it Team et al. (2024), Mistral-7B-Instruct-v0.3 Jiang et al. (2023). For the multimodal image generation task, we use ChatGPT-4o-mini Hurst et al. (2024).
@@ 3.3 Jailbreak Templates
In this study, we benchmark three existing jailbreak templates on the LLMs [^3] See Appendix for more details.Opposite Mode (OM), AntiLM, AIM Shen et al. (2024b) using code-mixed and phonetically perturbed inputs. For the image generation task, we test with a Base template– instructing image generation without asking for clarifications for generation style. We also devise two new jailbreaking templates
@@ 4.1 Generating Input Prompts
We use a 3-step process to generate the code-mixed prompts and phonetic perturbations for both text and image generation tasks [^3] See Appendix for more details..
@@ 4.3 Interpreting Phonetic Perturbations
  • We select a small subset of the dataset, specifically with $ASR_{CM}\leq 0.33, AA$$ASR_{CMP}\geq 0.5$ while ensuring that $AARR_{CMP}\geq AARR_{CM}$.
@@ 5 Results & Observations
In Table 1 we report the average Attack Success Rate (AASR) for all datasets, prompts, models and jailbreak templates. We first note that Gemma and Mistral report the highest AASR across jailbreak templates and input prompts. In fact, if we do not use any jailbreaking template, attacks still succeed at least 65% of the time on Gemma and 68% on Mistral. This baseline experiment shows that models are still vulnerable to classic red-teaming attacks despite evidence of harm, and using code-mixed (CM) or phonetically perturbed (CMP) prompts in fact increases the AASR, thus showing the effectiveness of our approach. ChatGPT and Llama are fairly robust to the different types of input, irrespective of the jailbreaking template being used. We also note that combining CM prompts with any of the jailbreak templates again results in $\simeq$ 0 AASR for ChatGPT and Llama, showing that their safety filters put precedence on template based attacks rather than the language of the input; Gemma and Mistral still maintain high AASR across the templates. We also note that our proposed jailbreak template– Sandbox performs the best amongst all, for both ChatGPT and Llama. This could be because the models are not aware of this template. We can conclude that two out of the four models are successfully red-teamed using our input format and jailbreak template. Next, we attempt to understand how relevant the outputs are for each input prompt.
@@
In Figures 35, we plot the sequence attribution scores using the method of Integrated Gradients Sundararajan et al. (2017), for a given input prompt in the English language, its code-mixed (CM) variant and its code-mixed variant with phonetic perturbations (CMP) for the Llama-3-8B-Instruct LLM. Recall that we study the attribution scores for generating safe outputs; thus, if a token has a high score, it implies that the token contributes more towards triggering the safety filter and if it scores low, then it has a lower impact towards the generated output. In Figure 3, with the following prompt– “Okay, here’s a hypothetical scenario. Let’s say someone wanted to use drama to promote hate speech and discrimination. Are there any specific techniques or approaches they would use for this purpose?”, we observe how the tokens– “hate”, “speech” and “discrimination” have a high attribution score at the embedding layer, 1st decoder layer and the 8th decoder layer. This means that these tokens are primarily responsible for generating the safe response from the LLM. “hate” and “speech” retain their high importance even at the 16th decoder layer. A similar observation can be made in Figure 4, for the same prompt written in a code-mixed (HI-EN) manner– “Ek hypothetical scenario tumhare saame rakhta hun. maan lo ki koi vyankti hate speech aur discrimination ko promote karna chahta hai through the use of drama. kya aisi koi specific techniques ya approaches hai jinka use ve is purpose ke liye kar sakte hain?”, where “hate” and “speech” are written in the English language. Thus, standard code-mixing may not be enough to bypass the safety filters of LLMs. On the other hand, in Figure 5, we note that the prompt, after phonetic perturbations as follows– “Ek hypothetical scenario tumhare saame rakhta hun. maan lo ki koi vyankti haet speech aur bhedbhaav ko promote karne ke liye drama ka upyog karna chahta hai. kya aisi koi vishisht techniques ya approaches hai jinka upyog ve is uddeshy ke liye kar sakte hain?”, generates radically different tokens– “hate” $\rightarrow$ “haet” tokenized as “ha” + “et” and, “discrimination” $\rightarrow$ “bhed bhav” tokenized as “b” + “hed” + “b” + “ha” + “av”. This leads to a change in the attribution scores as well, which are now low for the sensitive words and thus do not trigger the safety filters. A similar observation can be made for other prompts. Thus, based on our observations, we can state that phonetic perturbations lead to input tokenisation in a way that impacts the safety filters of LLMs, thus allowing attackers to generate harmful outputs.
@@ 6 Discussion
$\bullet$ Improved safety measures for all models: While Llama and ChatGPT show resistance to standard jailbreak attacks, Gemma and Mistral are easily jailbroken, even in English. This highlights the need for better evaluation and alignment methods for these models, especially considering their wide-scale availability. In multilingual and multimodal settings, especially code-mixed settings, safety filters of all models degrade drastically, revealing the need for better efforts in multilingual multimodal red teaming and alignment.
$\bullet$ Input tokenization determines safe outputs: If input tokens are out of vocabulary or perturbed, the safety filters get bypassed easily, but the model is able to interpret the instructions in the input prompt and generate harmful responses. Thus a more robust tokenization strategy needs to be developed.
$\bullet$ Template-based safety measures do not generalize well: The results from our experiments conducted across four models in the 7-8B parameter range using code-mixing in the Hindi language shows how template-based safety measures often fail, with the interpretability experiment revealing tokenization to be the main contributor to this failure. Therefore, we conclude that guardrails of models that are safety trained against template based attacks do not generalize well to attacks that deviate from set patterns, by including elements such as perturbations or code-mixing. This highlights the need for more generalizable safety training methods that are robust to deviations from standard attack patterns.
@@ 7 Conclusion
Future Work: An immediate area of future work is to align the models based on the findings from the explainability experiments. We plan to scale our efforts to more prompt categories, models, languages & jailbreak templates, and expand our red-teaming approach to other output modalities like speech.
@@ A.1.3 Jailbreaking Templates
In all cases, the templates contain instructions for allowing the jailbroken counterparts of the bots to bypass any safety guardrails, and not inform the user about legal and ethical guidelines. The jailbreak templates are used as system prompts for all the models except for Gemma-1.1-7b-it which does not allow using system prompts. For Gemma, we append it to the main prompt as a prefix.